OpenStudy (siberman):

Hello, just got this question wrorg apparently, and would love to know why. It is: simplify: log(5/36) - log(5/9) i did: log5 - log4 + 2log3/log5 - 2log3 = -log4/-1 = log4 i got the -1 on the bottom from cancelling 2log3/-2log3 Any explanation would be great, thanks heaps.

OpenStudy (campbell_st):

rewrite the equation using log laws \[\log( \frac{\frac{5}{36}}{\frac{5}{9}})\] divide by a fraction invert and multiply \[\log(\frac{5}{36} \times \frac{9}{5})\] this gives \[\log(\frac{1}{4})\] use the log law for division log(1) - log(4) = -log(4) since log(1) = 0

OpenStudy (campbell_st):

in your question you have log(5) - log(36) - log(5) + log(9) -2log(6) + 2 log(3) = -2log(2) - 2log(3) + 2log(3) please note = -2log(6) = 2(log(2 x 3)) = -2log(2) - 2log(3) to the answer is -2log(2) or -log(2^2) or -log(4)